# New binary black hole mergers in the LIGO–Virgo O3b data

### v2

@@ I Introduction
The Advanced LIGO and J Aasi et al. (2015) and Advanced Virgo Acernese et al. (2015) detectors have now been joined by KAGRA Akutsu et al. (2021) to form the LIGO–Virgo–KAGRA (LVK) collaboration, which recently released data from the second half of the third observing run (O3b) to the public Vallisneri et al. (2015). The LVK detected gravitational waves (GW) from 35 compact binary coalescence (CBC) events including two probable neutron star–black hole (NSBH) mergers Abbott et al. (2021). 32 of these were found to be coincident in the Hanford–Livingston (HL) detectors, among which 30 were reported to be binary black hole (BBH) signals. Nitz et al. (2021) independently produced the 4-OGC catalog, which included an additional 7 BBH mergers that were not in the LVK catalog.
In this work, we report the detection of 6 new candidate BBH merger signals from our coincident search of O3b HL data (excluding Virgo). Our search was carried out with a pipeline that is very similar to the one used in Olsen et al. (2022) (see Venumadhav et al. (2019) for an overview of the pipeline). We show these detections in Fig. 1 alongside the 90 CBC signals detected by LVK, additional events detected in 4-OGC, and the events previously detected by our pipeline in O1-O3a with probability of astrophysical origin () greater than one half Abbott et al. (2016a, 2019); Nitz et al. (2019, 2020); Abbott et al. (2021a); The LIGO Scientific Collaboration et al. (2021); Venumadhav et al. (2019, 2020); Olsen et al. (2022).
@@
Our pipeline confirmed the detections of 17 out of the 30 HL coincident BBH mergers reported in the GWTC-3 catalog (thus, in total, we detected 23 candidate CBC events in our coincident search of HL data from O3b). When we investigated our pipeline’s performance on the remaining events in the LVK catalog, we found that 6 of them were marginal in our catalog, i.e., recovered with a lower than one half, while 6 were vetoed for failing one or more of our signal consistency and excess power tests, and one event was missed. Most of the marginally recovered events were detected by only one of the LVK search pipelines, while three of the vetoed events show signs of physical effects like spin-precession and higher harmonics that were not modeled in our template bank. We also note that four of the vetoed events were flagged as having data-quality issues in the GWTC-3 catalog paper. Finally, we also checked that 4 of the 6 new candidates reported in this work were also present in the list of GWTC-3 sub-threshold candidates LIGO Scientific Collaboration and Collaboration (2021).
@@
We find that for at least two of the new events, detailed parameter estimation shows some evidence for effects such as higher-order modes (HM) and spin-precession, and also some information from the VIRGO data, which suggests that their significance could be potentially improved by including these elements in future improved searches (our current search of the Hanford–Livingston coincident data uses aligned-spin templates that only have the fundamental () harmonic). As noted earlier, some loud events in the LVK catalog that were vetoed by our pipeline also showed evidence for these physical effects, which additionally motivates the development of improved pipelines incorporating them.
@@ III.2 Low-mass sources
This event, with , has a bi-modal posterior distribution in the masses and effective spin under our default spin prior (see Fig. 12). The first posterior mode occurs at the secondary mass, and effective spin, while the second one occurs at and . Interestingly, the likelihoods of these two modes are very similar, which suggests that the data seems to prefer both modes equally likely. Both these solutions are, however, very interesting astrophysically. While the first solution hints at a highly negatively spinning stellar mass BBH merger, the second solution suggests at a significantly asymmetric () non-spining BBH merger with the secondary mass confidently lying in the LMG.
@@ IV Comparison to previously reported catalogs
The events GW191109_010717 and GW200129_065458 had fairly high SNR and are high-mass binaries with imprints of precession Abbott et al. (2021). Our search used a template bank that was built using the non-precessing approximant, and the signal consistency tests compare the data only to the best-fit waveform within the template bank – the absence of precession and higher-harmonics in the template can make it more likely for events that exhibit these phenomena to trigger the signal-consistency tests 333We note that some of the vetoed events here are restored in our search with higher-harmonics Wadekar et al. (2023b).. The data for two events was also flagged for data-quality issues in the GWTC3 catalog.
Our pipeline vetoed GW191113_071753 since the residuals after subtracting the best-fit template exhibited excess power beyond the expectation from Gaussian noise in multiple frequency bands. In the GWTC3 analysis, this candidate was marked as affected by transient, non-Gaussian detector noise. The best-fit waveform in PE for this candidate has a significantly asymmetric mass ratio , and the absence of higher modes in our template bank could have also contributed to the failure of the excess power tests.
@@
The other two events were straightforwardly vetoed in our search; GW191127_050227 failed multiple excess power tests in a few specific bands in a unit-second interval, while GW191103_012549 failed the split tests. We also note that the data for GW191127_050227 was flagged as being affected by data-quality issues in the GWTC3 catalog paper.
@@ V.1 Astrophysical implications of the new events
The evolution of massive stars can go through pair-instability phases where the stars are completely disrupted and no remnant is left behind. This produces a gap in the predicted distribution of BH mass. The exact boundary of such a gap is not well determined, as it depends on multiple uncertain factors such as the rate of carbon-to-oxygen burning in the core , angular momentum transport, metallicity, wind loss, etc. Marchant and Moriya (2020); Farmer et al. (2019); Mehta et al. (2022).
@@
The GW detectors are slightly biased towards positively aligned () BBH systems due to selection effects.888This is because the systems are relatively louder for fixed mass and distance. This means that we can observe these systems from a larger volume of space. Thus, their fraction in the population of GW observations is expected to be correspondingly larger than their fraction of the true astrophysical population. This is consistent with what we see in Fig. 1. This selection effect is indeed taken into consideration during the population analysis, and thus the inference of astrophysical parameters is not expected to be biased. However, it is not clear that the observation in Fig. 1 is entirely due to selection effects. We know that the anti-aligned systems are shorter in length and hence they could be more prone to being mistaken for glitches. The waveform approximants in this regime might also not be very faithful. One way to address these questions would be to do injection studies where we inject astrophysical signals into the data and let the pipeline discover them. We plan to do this in our next work.
@@ Appendix A Vetoes and Estimation of
where denotes the signal hypothesis and the noise hypothesis. Then, and denote the foreground and background trigger densities at the score , respectively. We compute following the same procedure outlined in Appendix B of Olsen et al. (2022). We measure the foreground and background trigger distributions directly from the data. To estimate the background, we generate an additional 2000 O3b runtimes worth of noise realizations from the O3b data using timeslides, i.e., unphysical time shifts between the data at the Hanford and Livingston detectors.
![Refer to caption](x4.png)
@@ Appendix B Posteriors for the new events
![Refer to caption](x5.png)
![Refer to caption](x6.png)
![Refer to caption](x7.png)
![Refer to caption](x8.png)
![Refer to caption](x9.png)
![Refer to caption](x10.png)
@@ Appendix C Bi-modality in posteriors: GW191228_085854 and the LVK analog GW200210_092254
![Refer to caption](x11.png)
![Refer to caption](x12.png)
![Refer to caption](x13.png)

### v3

@@ I Introduction
The Advanced LIGO and J Aasi et al. (2015) and Advanced Virgo Acernese et al. (2015) detectors have now been joined by KAGRA Akutsu et al. (2021) to form the LIGO–Virgo–KAGRA (LVK) collaboration, which recently released data from the second half of the third observing run (O3b) to the public Vallisneri et al. (2015). The LVK detected gravitational waves (GW) from 35 compact binary coalescence (CBC) events including two probable neutron star–black hole (NSBH) mergers Abbott et al. (2021). 32 of these were found to be coincident in the Hanford–Livingston (HL) detectors, among which 30 were reported to be binary black hole (BBH) signals. Nitz et al. (2021) independently produced the 4-OGC catalog, which included an additional 7 BBH mergers that were not in the LVK catalog. However, it also missed 8 events reported in GWTC-3 that have a probability of astrophysical origin () greater than 0.5. This emphasizes that different pipelines identify events near the threshold with different significance.
In this work, we perform a search analysis on the O3b HL coincident data (excluding Virgo) using a pipeline closely resembling the one used in Olsen et al. (2022) (see Venumadhav et al. (2019) for an overview of the pipeline). We report detections of a total of 23 compact binary mergers, 6 of which are reported for the first time here. The search did not recover 13 events reported in GWTC-3 and the additional 7 events reported in 4-OGC (more details follow later in the introduction, with a detailed description provided in section IV). We show these detections in Fig. 1 alongside the 90 CBC signals detected by LVK, additional events detected in 4-OGC, and the events previously detected by our pipeline in O1-O3a with greater than one half Abbott et al. (2016a, 2019); Nitz et al. (2019, 2020); Abbott et al. (2021a); The LIGO Scientific Collaboration et al. (2021); Venumadhav et al. (2019, 2020); Olsen et al. (2022).
@@
Our pipeline confirmed the detections of 17 out of the 30 HL coincident BBH mergers reported in the GWTC-3 catalog. When we investigated our pipeline’s performance on the remaining events in the LVK catalog, we found that 6 of them were marginal in our catalog, i.e., recovered with a lower than one half, while 6 were vetoed for failing one or more of our signal consistency and excess power tests, and one event was missed. Most of the marginally recovered events were detected by only one of the LVK search pipelines, while three of the vetoed events show signs of physical effects like spin-precession and higher harmonics that were not modeled in our template bank. We also note that four of the vetoed events were flagged as having data-quality issues in the GWTC-3 catalog paper. Finally, we also checked that 4 of the 6 new candidates reported in this work were also present in the list of GWTC-3 sub-threshold candidates LIGO Scientific Collaboration and Collaboration (2021).
@@
We find that for at least two of the new events, detailed parameter estimation shows some evidence for effects such as higher-order modes (HM) and spin-precession, and also some information from the VIRGO data, which suggests that their significance could be potentially improved by including these elements in future improved searches (our current search of the Hanford–Livingston coincident data uses aligned-spin templates that only have the fundamental () harmonic).
As noted earlier, some loud events in the LVK catalog that were vetoed by our pipeline also showed evidence for these physical effects, which additionally motivates the development of improved pipelines incorporating them. We note that the LVK catalog was also produced using aligned-spin templates limited to the harmonic, as in our case. However, our veto tests are comparatively more stringent, making it crucial for us to search for signals showing hints of HM or precession using templates that explicitly include these effects.
@@ III.2 Low-mass sources
This event, with , has a bi-modal posterior distribution in the masses and effective spin under our default spin prior (see Fig. 13). The first posterior mode occurs at the secondary mass, and effective spin, while the second one occurs at and . Interestingly, the likelihoods of these two modes are very similar, which suggests that the data seems to prefer both modes equally. Both these solutions are, however, very interesting astrophysically. While the first solution hints at a highly negatively spinning stellar mass BBH merger, the second solution suggests at a significantly asymmetric () non-spining BBH merger with the secondary mass confidently lying in the LMG.
@@ IV Comparison to previously reported catalogs
The events GW191109_010717 and GW200129_065458 had fairly high SNR and are high-mass binaries with imprints of precession Abbott et al. (2021). Our search used a template bank that was built using the non-precessing approximant, and the signal consistency tests compare the data only to the best-fit waveform within the template bank – the absence of precession and higher-harmonics in the template can make it more likely for events that exhibit these phenomena to trigger the signal-consistency tests 333We note that some of the vetoed events here are restored in our search with higher-harmonics Wadekar et al. (2023b).. The data for two events was also flagged for data-quality issues in the GWTC-3 catalog. Currently, the triggers that fail any of our separate veto tests are completely excluded from our catalog. In the future, we plan to devise a scheme for including the failed-test candidates back into our catalog, but with an appropriate penalty. This approach would ensure that high-significance events are not entirely removed due to the absence of certain physical effects in our templates, though this may come at the cost of slightly increasing the background in our search.
Our pipeline vetoed GW191113_071753 since the residuals after subtracting the best-fit template exhibited excess power beyond the expectation from Gaussian noise in multiple frequency bands. In the GWTC-3 analysis, this candidate was marked as affected by transient, non-Gaussian detector noise. The best-fit waveform in PE for this candidate has a significantly asymmetric mass ratio , and the absence of higher modes in our template bank could have also contributed to the failure of the excess power tests.
@@
The other two events were straightforwardly vetoed in our search; GW191127_050227 failed multiple excess power tests in a few specific bands in a unit-second interval, while GW191103_012549 failed the split tests. We also note that the data for GW191127_050227 was flagged as being affected by data-quality issues in the GWTC-3 catalog paper.
@@ V.1 Astrophysical implications of the new events
The evolution of massive stars can involve phases such as PPISN and PISN, where the stars either lose a significant fraction of their mass or are completely disrupted, leaving no remnant. As a result, PISN create a gap in the BH mass distribution, while PPISN can populate the mass spectrum below this gap. The precise boundaries of this gap remain uncertain, as they depend on several factors, including the rate of carbon-to-oxygen burning in the core (), angular momentum transport, metallicity, and wind loss. Marchant and Moriya (2020); Farmer et al. (2019); Mehta et al. (2022).
@@
The GW detectors are slightly biased towards positively aligned () BBH systems due to selection effects.888This is because the systems are relatively louder for fixed mass and distance. This means that we can observe these systems from a larger volume of space. Thus, their fraction in the population of GW observations is expected to be correspondingly larger than their fraction of the true astrophysical population. This is consistent with what we see in Fig. 1. This selection effect is indeed taken into consideration during the population analysis, and thus the inference of astrophysical parameters is not expected to be biased.
@@ Appendix A Vetoes and Estimation of
where denotes the signal hypothesis and the noise hypothesis. Then, and denote the foreground and background trigger densities at the score , respectively. We compute following the same procedure outlined in Appendix B of Olsen et al. (2022). We measure the foreground and background trigger distributions directly from the data. We show these distributions in Fig. 5. To estimate the background, we generate an additional 2000 O3b runtimes worth of noise realizations from the O3b data using timeslides, i.e., unphysical time shifts between the data at the Hanford and Livingston detectors.
In Fig. 4 we show as a function of the ranking score . Using the same fitting ansatz as Olsen et al. (2022), we find that
provides a good fit over the relevant range. Notice that the fit coefficients are different from the ones in Olsen et al. (2022), as we could have expected.